OpenStudy (anonymous):

1) Express (-2x^2-7x-13)/(x^3+4x-3(x^2)-12) in partial fractions. 2) Variables x and y are related by the equation y(e^a)=x^(k-1) where a and k are constants. When the graph of lnxy and lnx is drawn a straight line is obtained. Given that the line passes through (-2,11) and (1,2), find the value of a and k. (Is a=-5 and k=-3?) Thanks!

zepdrix (zepdrix):

\[\Large \frac{-2x^2-7x-13}{x^3-3x^2+4x-12}\]Hmmmmmmmmm were you able to factor it? I'm trying to remember how to factor a cube -_-

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

u can use splitting the mid term

OpenStudy (anonymous):

or else synthetic division to factor cubivc polynomials

OpenStudy (anonymous):

kkk

OpenStudy (anonymous):

i kind of tried factorising and manipulating it and i got 2+((4x^2-15x+11)/(x^3-3x^2+4x-12))

OpenStudy (anonymous):

kkk

OpenStudy (anonymous):

then group and factorise it

OpenStudy (anonymous):

im kinda stuck after this step cause i i do not know how to make it into a partial fraction after this>.<

OpenStudy (dumbcow):

factor by grouping \[(x^{3} -3x^{2})+(4x-12) = x^{2}(x-3) +4(x-3) = (x^{2}+4)(x-3)\] \[\frac{-2x^{2}-7x-13}{(x^{2}+4)(x-3)} = \frac{Ax +B}{x^{2}+4}+\frac{C}{x-3}\]

OpenStudy (dumbcow):

as far as 2nd part of post, you are correct \[\ln (xy) = k \ln x -a\] the line going through 2 points is: y = -3x +5 thus k = -3 a = -5